# An improvement toward global boundedness in a fully parabolic chemotaxis with singular sensitivity in any dimension

### v1

@@ Abstract
We are concerned with the following chemotaxis system
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\frac{u}{v}\nabla v\right),\\ v_{t}=\Delta v-u+v,\end{cases}$
in an open bounded domain with smooth boundary $\Omega\subset\mathbb{R}^{n}$ with $n\geq 3$. It is known that solutions are globally bounded when $0<\chi<\sqrt{\frac{2}{n}}$. In this paper, by considering the following energy functional
$F_{\lambda}(u,v)=\int_{\Omega}\frac{u^{p}}{v^{r}}+\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\int_{\Omega}v^{p-r}$
for $p>1$, $r=\frac{p-1}{2}$, and $\lambda>0$, we improve this result by showing that there exists $\chi_{0}>\sqrt{\frac{2}{n}}$ such that for any $\chi\in(0,\chi_{0})$, solutions exist globally and remain bounded.
@@ 1 Introduction
We consider the following system arising from a chemotaxis model in an open bounded domain $\Omega\subset\mathbb{R}^{n}$, where $n\geq 2$:
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\frac{u}{v}\nabla v\right),\qquad&\text{in }\Omega\times(0,T_{\rm max}),\\ v_{t}=\Delta v-v+u,\qquad&\text{in }\Omega\times(0,T_{\rm max}),\end{cases}$ (1.1)
@@
$\begin{cases}u_{0}\in C^{0}(\bar{\Omega})\text{ is nonnegative with }\int_{\Omega}u_{0}>0,\\ v_{0}\in W^{1,\infty}(\Omega),\quad v_{0}>0\quad\text{in }\bar{\Omega}.\end{cases}$ (1.3)
@@
The system (1.1) was originally introduced in the seminal work [7], where it was proposed as a mathematical model to describe the phenomenon of chemotaxis—namely, the directed movement of cells or microorganisms toward or away from chemical signals in their environment. This mechanism plays a fundamental role in a variety of biological processes, such as the aggregation of cellular slime molds, immune response behavior, and the formation of patterns in bacterial populations. In the context of system (1.1), the function $u(x,t)$ typically denotes the density of a cell population (such as bacteria or amoebae), while $v(x,t)$ represents the concentration of a chemical substance that influences the movement of these cells at a given spatial location $x$ and time $t$. This interaction reflects the phenomenon of chemotaxis, where cellular motion is guided by gradients of chemical signals. A well-studied variant of this model arises when the second equation in system (1.1) is replaced by the elliptic equation $0=\Delta v-v+u$, leading to the so-called parabolic-elliptic system. Under the assumption of radially symmetric initial data, it was shown in [17] that solutions to this system exist globally in time whenever the chemotactic sensitivity parameter satisfies $\chi<\frac{2}{n-2}$, whereas solutions may blow up if $\chi>\frac{2n}{n-2}$. This result was later extended in [2], where the condition for global existence was relaxed to $\chi<\frac{2}{n}$ without requiring radial symmetry in the initial data. Moreover, global boundedness of solutions under this condition was established in [6]. In two spatial dimensions, a stronger result was obtained in [5], demonstrating that blow-up does not occur for any positive value of $\chi$, highlighting a critical difference between low and high-dimensional behavior. More recently, in [8], it was shown that there exists a threshold $\chi_{0}>\frac{2}{n}$ such that for all $\chi<\chi_{0}$, solutions are globally defined and remain uniformly bounded for all time. Despite these substantial contributions, the exact critical value of $\chi$ that delineates the transition between global existence and finite-time blow-up for the parabolic-elliptic system is still not fully characterized. For further advances related to parabolic-elliptic systems with singular sensitivity, we refer the reader to [23, 24, 13, 10, 9, 14].
We now turn our attention to the fully parabolic Keller–Segel system. It was established in [21] that classical solutions exist globally whenever $\chi<\sqrt{\frac{2}{n}}$, and later in [3] that such solutions remain uniformly bounded under the same condition. Besides, in [21], it was shown that global weak solutions exist when $\chi<\sqrt{\frac{n+2}{3n-4}}$. By defining a slightly more general concept of weak solutions, the authors in [18] proved the global existence of radial solutions when $\chi<\sqrt{\frac{n}{n-2}}$. Later, the construction of global generalized solutions was established in [12] under the restriction that
$\chi<\begin{cases}\infty\qquad\text{if }n=2,\\ \sqrt{8}\qquad\text{if }n=3,\\ \frac{n}{n-2}\qquad\text{if }n\geq 4.\end{cases}$
Therefore, these results naturally lead to the question of whether there exists a critical value $\chi_{c}>0$ such that solutions are globally bounded for all $\chi<\chi_{c}$, whereas for $\chi>\chi_{c}$ blow-up may occur. The result in [11] suggests that this critical threshold satisfies $\chi_{c}>1.016$ when $n=2$, indicating that $\chi=1$ is not the critical value in dimension two. Further, the authors in [4] conjectured that $\chi_{c}=\frac{n}{n-2}$, and provided a rigorous proof that solutions are global and bounded whenever $\chi<\frac{n}{n-2}$, assuming the left hand side of the first equation in system (1.1) is replaced by $\tau u_{t}$ with sufficiently small $\tau>0$. Without imposing smallness on $\tau$, the recent work [1] improved previous lower bound for $\chi_{c}$ in low dimensions by showing $\chi_{c}\geq\frac{4}{n}$. This result enhances those in [21] and [11] for dimensions $2\leq n\leq 8$, since $\frac{4}{n}\geq\sqrt{\frac{2}{n}}$ and $2>1.016$. In line with the result established in [4], it was shown in [22] that solutions are arbitrary large if $\tau$ is sufficiently small and $\chi>\frac{n}{n-2}$. This result clearly supports the aforementioned conjecture that $\chi_{c}=\frac{n}{n-2}$. Therefore, it is reasonable to conclude that the threshold $\chi_{c}$ exceeds $\sqrt{\frac{2}{n}}$, indicating that $\sqrt{\frac{2}{n}}$ is not the critical value in any dimension. This conclusion has been rigorously established for spatial dimensions up to eight or when $\tau$ is sufficiently small and the domain $\Omega$ is convex; however, whether it holds for dimensions $n\geq 9$, or without the smallness condition on $\tau$ and in nonconvex domains, remains an open problem. In this work, we address this gap by proving that $\chi_{c}>\sqrt{\frac{2}{n}}$ for all $n\geq 3$. Our main theorem is stated as follows:
@@
Let $\Omega\subset\mathbb{R}^{n}$, with $n\geq 3$, be an open bounded domain with smooth boundary. Then there exists $\chi_{0}>\sqrt{\frac{2}{n}}$ such that, for any $\chi\in(0,\chi_{0})$, the system (1.1), with the boundary condition (1.2) and initial condition (1.3), admits a unique global positive classical solution $(u,v)\in\left[C^{2,1}\left(\bar{\Omega}\times(0,\infty)\right)\cap C^{0}\left(\bar{\Omega}\times[0,\infty)\right)\right]^{2}$ such that
@@
An explicit lower bound for $\chi_{0}$, established in the proof of Lemma 4.1, is $\sqrt{\frac{2(1+\delta)}{n}}$ with $\delta>\frac{(n-2)^{2}}{256(n-1)\left(16^{n}(3n^{2})^{n+1}+(912n^{2})^{n+1}\right)}$.
@@
The works [11], [4], and [1] all require the convexity of the domain $\Omega$, whereas Theorem 1.1 imposes no such condition.
In addition, we propose a novel approach for studying chemotaxis systems with singular sensitivity in arbitrary dimensions, drawing inspiration from the works [21], [11], and [19]. In [21], the global existence of solutions for $\chi<\sqrt{\frac{2}{n}}$ was established by analyzing the quantity $\int_{\Omega}u^{p}v^{-r}$ for $p>1$ and $r>0$. Subsequently, [11] suggested that the inclusion of an additional summand could yield improved results. In a different setting, [19] introduced an energy functional involving the term $\int_{\Omega}v^{-q+1}|\nabla v|^{q}$, which motivates us to consider the generalized form $\int_{\Omega}v^{-l}|\nabla v|^{2p}$. This leads to the identification of $\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$ as a suitable additional term to the original expression $\int_{\Omega}u^{p}v^{-r}$. To further facilitate the analysis, we introduce the auxiliary term $\int_{\Omega}v^{p-r}$, which allows us to directly estimate the boundedness of $\int_{\Omega}u^{p}v^{-r}$ without relying on the iterative argument employed in [3, 16] and without requiring a prior bound on $\int_{\Omega}v^{p-r}$. Consequently, we construct the energy functional
$F_{\lambda}(u,v)=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\int_{\Omega}v^{p-r},\quad\text{with }r=\frac{p-1}{2},$
which enables us to establish an alternative proof of the global boundedness of solutions to (1.1). In particular, it can be shown—using Lemma 3.1—that $F_{0}(u,v)$ remains bounded whenever $\chi<\sqrt{\frac{2}{n}}$.
@@ 2 Preliminaries
Let us begin this section by recalling the local existence of and an extensibility criteria of solutions to the system (1.1) as established in [11][Theorem 2.3].
@@
$u$ $\in C^{0}\left(\bar{\Omega}\times[0,T_{\rm max})\right)\cap C^{2,1}\left(\bar{\Omega}\times(0,T_{\rm max})\right)\quad\text{and }$
@@
From now on, we denote $(u,v)$ as a classical solution to (1.1) in $(0,T_{\rm max})$. The following lemma gives an uniform in time $L^{1}$ bound for $v$.
@@
$\int_{\Omega}u(\cdot,t)=m\qquad\text{for all }t\in(0,T_{\rm max}),$
@@
$\int_{\Omega}v(\cdot,t)\leq C\qquad\text{for all }t\in(0,T_{\rm max}).$
@@
$\frac{d}{dt}\int_{\Omega}u(\cdot,t)=0\qquad\text{for all }t\in(0,T_{\rm max}),$
@@
$\int_{\Omega}u(\cdot,t)=m:=\int_{\Omega}u_{0}\qquad\text{for all }t\in(0,T_{\rm max}).$
@@
$\frac{d}{dt}\int_{\Omega}v+\int_{\Omega}v=m\qquad\text{for all }t\in(0,T_{\rm max}).$
Applying Gronwall’s inequality to this implies that
@@
$\inf_{x\in\Omega}v(x,t)\geq\eta\qquad\text{for all }t\in(0,T_{\rm max}).$
@@
    $\left\|v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\quad\text{for all }t\in(0,T).$
@@
    $\left\|\nabla v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\quad\text{for all }t\in(0,T).$
@@
and employing integration by parts yields that
@@
which further implies (2.2). We use integration by parts and Hölder’s inequality to derive (2.3) by estimating
@@
$\quad-\int_{\Omega}\phi|\nabla\ln\phi|^{l+1}\nabla\ln\phi\cdot\nabla|\nabla\phi|^{q-l-1}$
@@
$\quad-(q-l-1)\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\phi\cdot\nabla\phi$
@@
Combining (2) and (2.6) with straightforward simplifications deduces that
@@
Dropping the second term of the right hand side and using Young’s inequality entails that
@@
which deduces (2.4). The proof is now complete.
@@
$\frac{\partial|\nabla\phi|^{2}}{\partial\nu}\leq c_{1}|\nabla\phi|^{2}\quad\text{on }\partial\Omega,$
@@
$\int_{\partial\Omega}|w|\leq c_{2}\int_{\Omega}|\nabla w|+c_{2}\int_{\Omega}|w|\quad\text{for all }w\in C^{1}(\bar{\Omega}).$
@@
$\leq c_{1}c_{2}\int_{\Omega}\left|\phi^{-p-r}|\nabla\phi|^{2p}\right|+c_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p}.$ (2.10)
Using the identity $\nabla|\nabla\phi|^{2}=2D^{2}\phi\cdot\nabla\phi$ and Young’s inequality implies that
$c_{1}c_{2}\int_{\Omega}\left|\phi^{-p-r}|\nabla\phi|^{2p}\right|$ $\leq 2pc_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-1}|D^{2}\phi|+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}$
@@
$\quad+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}.$ (2.11)
Abbreviating $c_{3}:=c_{1}c_{2}+\frac{p^{2}c_{1}^{2}c_{2}^{2}}{\epsilon}$ and $c_{4}:=c_{1}c_{2}(p+r)$, and applying Young’s inequality with $\epsilon>0$ deduces that
@@
Combining estimates from (2) to (2) entails that
@@
where $c_{5}=\frac{2^{p}}{\epsilon^{p}}c_{3}+\left(\frac{2}{\epsilon}\right)^{2p+1}c_{4}^{2p+2}$. Thanks to Lemma 2.5, one can find $c_{6}=c_{6}(p,r,n)>0$ such that
$\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}|D^{2}\phi|^{2}+\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}\leq c_{6}\int_{\Omega}\phi^{-p-r+2}|\nabla\phi|^{2p-2}|D^{2}\ln\phi|^{2}$

### v2

@@ Abstract
This paper deals with the problem of global solvability and boundedness of classical solutions to a fully parabolic chemotaxis system with singular sensitivity in any dimensional setting. In particular, We show that the system
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\dfrac{u}{v}\nabla v\right),\\ v_{t}=\Delta v-v+u,\end{cases}$
posed in a bounded domain $\Omega\subset\mathbb{R}^{n}$ with $n\geq 3$, admits a global bounded classical solution provided that $\chi\in(0,\chi_{0})$ with $\chi_{0}>\sqrt{\frac{2}{n}}$ can be determined explicitly. This result extends several existing works, which established global boundedness under the more restrictive condition $\chi<\sqrt{\frac{2}{n}}$, and shows that this threshold is not an optimal upper bound for preventing blow-up.
@@ 1 Introduction
We consider the following system arising from a chemotaxis model in a bounded domain $\Omega\subset\mathbb{R}^{n}$, where $n\geq 3$:
$\begin{cases}u_{t}=\Delta u-\chi\nabla\cdot\left(\frac{u}{v}\nabla v\right),\hskip 18.49988pt&\text{in }\Omega\times(0,T_{\rm max}),\\ v_{t}=\Delta v-v+u,\hskip 18.49988pt&\text{in }\Omega\times(0,T_{\rm max}),\end{cases}$ (1.1)
@@
$\begin{cases}u_{0}\in C^{0}(\bar{\Omega})\text{ is nonnegative with }\int_{\Omega}u_{0}>0,\\ v_{0}\in W^{1,\infty}(\Omega),\qquad v_{0}>0\qquad\text{in }\bar{\Omega}.\end{cases}$ (1.3)
@@
The system (1.1) was originally introduced in the seminal work [8], where it was proposed as a mathematical model to describe the phenomenon of chemotaxis—namely, the directed movement of cells or microorganisms toward or away from chemical signals in their environment. This mechanism plays a fundamental role in a variety of biological processes, such as the aggregation of cellular slime molds, immune response behavior, and the formation of patterns in bacterial populations. In the context of system (1.1), the function $u(x,t)$ typically denotes the density of a cell population (such as bacteria or amoebae), while $v(x,t)$ represents the concentration of a chemical substance that influences the movement of these cells at a given spatial location $x$ and time $t$.
1.1 Literature review
In this subsection, let us briefly recall some mathematical results associated with the system (1.1).
A well-studied variant of system (1.1) arises when the second equation is replaced by the elliptic equation
$0=\Delta v-v+u,$
leading to the so-called parabolic–elliptic chemotaxis system. Under the assumption of radially symmetric initial data, it was shown in [23] that solutions exist globally in time provided that the chemotactic sensitivity satisfies $\chi<\frac{2}{n-2},$ whereas finite-time blow-up may occur if $\chi>\frac{2n}{n-2}$. This result was subsequently strengthened in [2], where it was shown that the condition $\chi<\frac{2}{n}$ guarantees the global existence of solutions for arbitrary initial data. Moreover, global boundedness of solutions under this assumption was established in [7].
In two spatial dimensions, a substantially stronger result was obtained in [5], where it was shown that blow-up is excluded for all $\chi>0$, thereby revealing a fundamental qualitative distinction between two-dimensional and higher-dimensional dynamics. Very recently, the author in [9] further improved these results in higher dimensional domains by showing that the parabolic-elliptic system (1.1) possesses a global bounded classical solution provided that
$\chi<\frac{2}{n}+\frac{2n-1}{2n^{3}}\sqrt{\frac{n}{2n+2}},\qquad n\geq 3.$
Despite these substantial advances, the precise critical value of $\chi$ separating global existence from finite-time blow-up in the parabolic–elliptic setting remains an open problem. For further developments concerning parabolic–elliptic chemotaxis systems with singular sensitivity, we refer the reader to [30, 31, 14, 11, 10, 18].
We now turn our attention to the fully parabolic case. In two spatial dimensions, it was shown that the condition $\chi\leq 1$ guarantees the global solvability of solutions to system (1.1) [2]. This threshold was later slightly improved, and blow-up was excluded whenever $\chi<1.015$ [12]. Moreover, under the additional assumption that the initial data are radially symmetric, global boundedness was obtained for a modified system in which the left-hand side of the second equation is replaced by $\tau v_{t}$, provided that $\tau>0$ is sufficiently small; in this setting, solutions remain global and bounded for arbitrary $\chi>0$ [6].
In higher-dimensional domains, global existence of solutions was established, and global boundedness was further derived under the condition
$\chi<\sqrt{\frac{2}{n}}$
[27, 3]. If the left-hand side of the first equation is replaced by $\tau u_{t}$ with sufficiently small $\tau>0$, then global existence and boundedness hold for all
$\chi<\frac{n}{n-2}$
[5]. In contrast, when $\tau$ is sufficiently small and $\chi>\frac{n}{n-2}$, solutions may become arbitrarily large [28]. More recently, without imposing any smallness assumption on $\tau$, it was proved in [1] that the condition
$\chi<\frac{4}{n}$
ensures global existence and boundedness of solutions, thereby improving earlier results [27, 12]. Beyond classical solutions, the existence of global weak solutions was obtained under the restriction
$\chi<\sqrt{\frac{n+2}{3n-4}}$
[27]. By introducing a slightly more general notion of weak solutions, global existence of radially symmetric solutions was further established provided that
$\chi<\sqrt{\frac{n}{n-2}}$
[24]. Finally, the existence of global generalized solutions was shown under the assumption
$\chi<\begin{cases}\infty,&\text{if }n=2,\\[2.77501pt] \sqrt{8},&\text{if }n=3,\\[2.77501pt] \dfrac{n}{n-2},&\text{if }n\geq 4,\end{cases}$
see [13]. For further developments concerning solvability and long-time behavior of variants of system (1.1), we refer the reader to [32, 29, 15, 16, 17, 19, 20].
1.2 Motivation and Contribution
In two spatial dimensions, the author in [2] investigated the following energy functional to obtain global existence of solutions to (1.1)
$A(t)=\int_{\Omega}u\ln u-a\int_{\Omega}u\ln v\qquad\text{for all }t>0,$
where $a>0$. This result was later refined in [12], where the author demonstrated that global existence and boundedness by employing an additional term $\int_{\Omega}\frac{|\nabla v|^{2}}{v}$ to the functional $A(t)$. More precisely, the functional
$B(t)=\int_{\Omega}u\ln u-a\int_{\Omega}u\ln v+b\int_{\Omega}\frac{|\nabla v|^{2}}{v}\qquad\text{for all }t>0,$
with $b>0$, was investigated to obtain a stronger threshold for $\chi$.
In higher dimensions ($n\geq 3$), as mentioned before, [27] established the existence of a global classical solution under the assumption $\chi<\sqrt{\frac{2}{n}}$. The proof relied on the functional
$C(t)=\int_{\Omega}u^{p}v^{-r}\qquad\text{for all }t>0,$
where $p>1$ and $r>1$. This naturally raises a question analogous to that posed in [12]: namely, whether the approach developed in [27] already yields a sharp criterion for global boundedness, or whether it can be refined in a manner similar to the two-dimensional case.
Motivated by the aforementioned results, we propose a new approach showing that the value $\sqrt{\frac{2}{n}}$ is not the sharp upper bound for blow-up prevention in arbitrary spatial dimensions. To this end, we introduce the energy functional
$F_{\lambda}(u,v)=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\int_{\Omega}v^{p-r}\qquad\text{for all }t>0,$
where $p>1$, $r=\frac{p-1}{2}$, and $\lambda>0$.
The inclusion of the additional term
$\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$
provides extra flexibility in the energy estimates and allows us to extend the admissible upper bound on $\chi$ beyond that obtained in [27]. The construction of this term is inspired by [25], where an energy functional containing
$\int_{\Omega}\frac{|\nabla v|^{q}}{v^{q-1}}$
was introduced to address chemotaxis-consumption systems with weakly singular sensitivity.
In addition, the presence of the term $\int_{\Omega}v^{p-r}$ in the functional $F_{\lambda}(u,v)$ enables us to derive a direct bound for $\int_{\Omega}u^{p}v^{-r}$, thereby avoiding the iterative arguments used in [3, 22] and eliminating the need for any a priori estimate on $\int_{\Omega}v^{p-r}$.
1.3 Main result
Our main result on the global existence and boundedness of solutions to system (1.1) can now be stated as follows.
@@ 1.3 Main result
Let $\Omega\subset\mathbb{R}^{n}$, with $n\geq 3$, be a bounded domain with smooth boundary. Then there exists $\chi_{0}>\sqrt{\frac{2}{n}}$ such that, for any $\chi\in(0,\chi_{0})$, the system (1.1), with the boundary condition (1.2) and initial condition (1.3), admits a unique global positive classical solution $(u,v)\in\left[C^{2,1}\left(\bar{\Omega}\times(0,\infty)\right)\cap C^{0}\left(\bar{\Omega}\times[0,\infty)\right)\right]^{2}$ such that
@@
An explicit lower bound for $\chi_{0}$ is established in Lemma 4.1, namely
$\chi_{0}\geq\sqrt{\frac{2(1+\delta)}{n}},$
where
$\delta=\frac{L_{2}L_{4}}{2L_{3}L_{5}},$
with
$L_{2}$ $=\frac{n-2}{16},\hskip 18.49988ptL_{3}=16^{n}\left(\frac{(2n-1)(3n-1)}{2}\right)^{n+1}+(304n(2n-2+\sqrt{n}))^{n+1},$
$L_{4}$ $=\frac{n-2}{8},\hskip 18.49988ptL_{5}=\frac{n-1}{4}.$
Moreover, one can verify that
$\delta>\frac{(n-2)^{2}}{256(n-1)\left(16^{n}(3n^{2})^{n+1}+(912n^{2})^{n+1}\right)}.$
@@
The works [12, 4, 1] require the convexity of the domain $\Omega$, whereas Theorem 1.1 imposes no such assumption. This is due to Lemma 2.6, which enables control of the nonlinear boundary integrals through the diffusion term without relying on the convexity of $\Omega$.
From a biological perspective, the result of Theorem 1.1 indicates that when chemotactic sensitivity is weak, the combined effects of diffusion and signal degradation are sufficient to regulate cell movement, thereby preventing excessive aggregation in any spatial dimension. This mechanism stands in sharp contrast to classical chemotaxis systems, where diffusion alone cannot balance chemotactic accumulation, often leading to pronounced clustering of the population.
Future work. An important open question is the optimal choice of $(r,\lambda)$ that maximizes the value of $\chi_{0}$, ensuring that for any $\chi<\chi_{0}$ the solutions exist globally and remain uniformly bounded in time. Another interesting direction for future research is to explore applications of the boundedness of
$\int_{\Omega}\frac{u^{p}}{v^{r}}+\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}$
in preventing blow-up for chemotaxis systems with (weak) singular sensitivity and logistic growth, or for chemotaxis-fluid models with singular sensitivity. In such models, as can be seen, a uniform lower bound for $v$ is generally no longer guaranteed due to the presence of logistic sources or fluid interactions. We leave these questions as topics for future investigations.
@@ 2 Preliminaries
Let us begin this section by recalling the local existence and an extensibility criterion of solutions to the system (1.1) as established in [12][Theorem 2.3].
@@
$u$ $\in C^{0}\left(\bar{\Omega}\times[0,T_{\rm max})\right)\cap C^{2,1}\left(\bar{\Omega}\times(0,T_{\rm max})\right)\qquad\text{and }$
@@
From now on, we denote $(u,v)$ as a classical solution to (1.1) in $(0,T_{\rm max})$. The following lemma gives a uniform in time $L^{1}$ bound for $v$.
@@
$\int_{\Omega}u(\cdot,t)=m\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$
@@
$\int_{\Omega}v(\cdot,t)\leq C\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
@@
$\frac{d}{dt}\int_{\Omega}u(\cdot,t)=0\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$
@@
$\int_{\Omega}u(\cdot,t)=m:=\int_{\Omega}u_{0}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
@@
$\frac{d}{dt}\int_{\Omega}v+\int_{\Omega}v=m\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
Applying Gronwall’s inequality to this, it follows that
@@
$\inf_{x\in\Omega}v(x,t)\geq\eta\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
@@
    $\left\|v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\qquad\text{for all }t\in(0,T).$
@@
    $\left\|\nabla v(\cdot,t)\right\|_{L^{\mu}(\Omega)}\leq C\left(1+\sup_{s\in(0,T)}\left\|u(\cdot,s)\right\|_{L^{\theta}(\Omega)}\right)\qquad\text{for all }t\in(0,T).$
@@
and employing integration by parts yields
@@
which further implies (2.2). We use integration by parts and Hölder’s inequality to derive (2.3)
@@
$\qquad-\int_{\Omega}\phi|\nabla\ln\phi|^{l+1}\nabla\ln\phi\cdot\nabla|\nabla\phi|^{q-l-1}$
@@
$\qquad-(q-l-1)\int_{\Omega}\phi^{-l-1}|\nabla\phi|^{q-2}\nabla\phi\cdot D^{2}\phi\cdot\nabla\phi$
@@
Combining (2) and (2.6) with straightforward simplifications, we deduce that
@@
Dropping the second term of the left hand side and using Young’s inequality, we derive
@@
which proves (2.4). The proof is now complete.
@@
$\frac{\partial|\nabla\phi|^{2}}{\partial\nu}\leq c_{1}|\nabla\phi|^{2}\qquad\text{on }\partial\Omega,$
@@
$\int_{\partial\Omega}|w|\leq c_{2}\int_{\Omega}|\nabla w|+c_{2}\int_{\Omega}|w|\qquad\text{for all }w\in C^{1}(\bar{\Omega}).$
@@
$\leq c_{1}c_{2}\int_{\Omega}\left|{\nabla}\left(\phi^{-p-r}|\nabla\phi|^{2p}\right)\right|+c_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p}.$ (2.10)
Using the identity $\nabla|\nabla\phi|^{2}=2D^{2}\phi\cdot\nabla\phi$ and Young’s inequality, it follows that
$c_{1}c_{2}\int_{\Omega}\left|\nabla\left(\phi^{-p-r}|\nabla\phi|^{2p}\right)\right|$ $\leq 2pc_{1}c_{2}\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-1}|D^{2}\phi|+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}$
@@
$\qquad+c_{1}c_{2}(p+r)\int_{\Omega}\phi^{-p-r-1}|\nabla\phi|^{2p+1}.$ (2.11)
Abbreviating $c_{3}:=c_{1}c_{2}+\frac{p^{2}c_{1}^{2}c_{2}^{2}}{\epsilon}$ and $c_{4}:=c_{1}c_{2}(p+r)$, and applying Young’s inequality with $\epsilon>0$, we derive that
@@
Combining estimates from (2) to (2) yields
@@
where $c_{5}=\frac{2^{p}}{\epsilon^{p}}c_{3}+\left(\frac{2}{\epsilon}\right)^{2p+1}c_{4}^{2p+2}$. Thanks to (2.3) and (2.4), one can find $c_{6}=c_{6}(p,r,n)>0$ such that
$\int_{\Omega}\phi^{-p-r}|\nabla\phi|^{2p-2}|D^{2}\phi|^{2}+\int_{\Omega}\phi^{-p-r-2}|\nabla\phi|^{2p+2}\leq c_{6}\int_{\Omega}\phi^{-p-r+2}|\nabla\phi|^{2p-2}|D^{2}\ln\phi|^{2}.$